OpenStudy (anonymous):

Find the sum of the first 30 terms of the sequence below. an=4n+1

OpenStudy (anonymous):

an=4(1)+1=5 then you do an=4(2)+1=9 then so on and so forth until it is an=4(30)+1=121 good luck

OpenStudy (anonymous):

you understand?

OpenStudy (anonymous):

there is a better way

OpenStudy (anonymous):

that will take forever

OpenStudy (ranga):

First term is: 4(1) + 1 Second term is: 4(2) + 1 Third term is: 4(3) + 1 ... ... 30th term is: 4(30) + 1 Add them all up: 4(1 + 2 + 3 + ..... + 30) + (1 + 1 + 1 ... 30 times) The first parenthesis is the sum of the first 30 natural numbers. There is a standard formula for that: n(n+1) / 2 where n = 30 The second parenthesis has the same number 1 repeating 30 times. Now you can addd them up.

OpenStudy (anonymous):

the 1 is repeated 30 time so the sum of number 1 is 30 and we know that the \[\sum_{1}^{ n} n = [n*(n+1)]/2\] so we have 4*[n*(n+1)/2]+30 and the n is 30

OpenStudy (anonymous):

understand??

OpenStudy (campbell_st):

the really easy method is \[S_{30} = \frac{30}{2}\times[5 + 121]\]

OpenStudy (campbell_st):

the sequence is arithmetic, the 1st term is 5 and the last term is 121. so add the 1st and last terms then multiply by 15, half the number of terms

OpenStudy (anonymous):

1890 is the answer?

OpenStudy (ranga):

Yes, 1890.

OpenStudy (campbell_st):

the formula for 1st and last terms is \[S_{n} = \frac{n}{2} \times (a + l)\] a = 1st term, l = last term and n = number of terms

OpenStudy (campbell_st):

yes, thats the answer, well done

OpenStudy (anonymous):

thank you